OpenStudy (anonymous):

Lisa invests $4,000 in two types of bonds, bond A and bond B. Bond A offers a 10% return, and bond B offers a 6% return. Lisa invests $x in bond A and $y in bond B. Her total return on the investment is $340. The system of linear equations defining the situation is x+y=4,000 and .1x=.06y=340. The amount Lisa invested at the rate of 10% is ___ , and the amount she invested at the rate of 6% is ___ .

OpenStudy (anonymous):

I think I got the linear equation right, Now how do I find how much she invested on each at 10% and 6%

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Hero

OpenStudy (amistre64):

Lisa invests $4,000 in two types of bonds, bond A and bond B. A + B = 4000 Bond A offers a 10% return, and bond B offers a 6% return. .10 A + .06 B = 340 Her total return on the investment is $340. solve the system of equations with whatever method strikes your fancy

OpenStudy (amistre64):

since the relationship A+B=4000 must satisfy the setup for all values A and B we can state that A = 2000 + n, and B = 2000 -n afterall: 2000 +n + 2000 - n = 4000 for all n .10(2000+n) + .06(2000-n) = 340, once we know n we know A and B

OpenStudy (anonymous):

I figured it out I forgot to close, I think it was 1500 for b or 2500 for a (: Thanks though